# ASPD: Unlocking Adaptive Serial-Parallel Decoding by Exploring Intrinsic Parallelism in LLMs

### v1

@@ Abstract
The increasing scale and complexity of large language models (LLMs) pose significant inference latency challenges, primarily due to their autoregressive decoding paradigm characterized by the sequential nature of next-token prediction. By re-examining the outputs of autoregressive models, we observed that some segments exhibit parallelizable structures, which we term intrinsic parallelism. Decoding each parallelizable branch simultaneously (i.e. parallel decoding) can significantly improve the overall inference speed of LLMs. In this paper, we propose an Adaptive Serial-Parallel Decoding (ASPD), which addresses two core challenges: automated construction of parallelizable data and efficient parallel decoding mechanism. More specifically, we introduce a non-invasive pipeline that automatically extracts and validates parallelizable structures from the responses of autoregressive models. To empower efficient adaptive serial-parallel decoding, we implement a Hybrid Decoding Engine which enables seamless transitions between serial and parallel decoding modes while maintaining a reusable KV cache, maximizing computational efficiency. Extensive evaluations across General Tasks, Retrieval-Augmented Generation, Mathematical Reasoning, demonstrate that ASPD achieves unprecedented performance in both effectiveness and efficiency. Notably, on Vicuna Bench, our method achieves up to 3.19x speedup (1.85x on average) while maintaining response quality within 1% difference compared to autoregressive models, realizing significant acceleration without compromising generation quality. Our framework sets a groundbreaking benchmark for efficient LLM parallel inference, paving the way for its deployment in latency-sensitive applications such as AI-powered customer service bots and answer retrieval engines.
@@ 1 Introduction
Figure 1: Overview of Data Intrinsic and Model Internal Parallelism. Proportion of Parallel Data (PPD). The ratio of parallelizable samples in the dataset; Degree of Parallelism (DP). The proportion of parallel tokens among all tokens in the output; Average Branch Number (ABN). The average number of parallel branches in the output.
Through extensive analysis of LLM outputs, we have made a crucial observation: despite their sequential generation pattern, a substantial portion of the model’s outputs exhibits inherent parallelism, which can be harnessed for concurrent processing. As shown in Figure 1, across various scenarios including general [^1] https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered, retrieval-augmented-generation [^2] https://huggingface.co/datasets/neural-bridge/rag-dataset-12000, mathematical reasoning [^3] https://huggingface.co/datasets/open-r1/OpenR1-Math-220k , and our internal multi-turn reading comprehension benchmarks, model responses consistently reveal substantial potential for parallelization. Within these parallelizable components, different scenarios exhibit varying levels of parallelism, ranging from 30.6% (Math) to 70.2% (General). By harnessing these inherently parallelizable responses for concurrent output, the model’s response speed can be significantly enhanced.
@@
Recent attempts to address these challenges have yielded promising but limited solutions. SoT (Ning et al., 2023) introduced a two-stage approach using response skeletons and parallel generation through external APIs, but this method incurs additional latency due to KV-cache limitations and increased memory overhead with batch processing. APAR (Liu et al., 2024) proposed discarding parallel branch KV-caches during final generation, which, while addressing position encoding issues, compromises response quality. PASTA (Jin et al., 2025) explored asynchronous parallel decoding with pre-allocated position ranges, but struggles with position encoding mismatches when actual generation lengths deviate from predictions.
@@
1) Non-invasive parallel data transformation pipeline. We develop an innovative pipeline that automatically discovers and extracts inherent parallelizable structures from autoregressive model responses, which identifies semantically independent components that can be processed concurrently while preserving the response’s original style. This enables us to build high-quality parallel training corpora without altering the response probability distribution or requiring manual annotation.
2) Internal parallelization module. Our framework introduces a custom parallelization mask that enables efficient parallel processing without the overhead of batching or threading, maintaining consistency with traditional autoregressive decoding patterns.
3) Hybrid Decoding Engine support efficient and cyclic serial-parallel decoding. We propose novel parallel positional encoding and training techniques that enable zero-overhead integration of parallel branches and support multiple rounds of parallel decoding.
4) Comprehensive evaluation achieving optimal balance between effectiveness and efficiency. Through extensive evaluation across diverse benchmarks - including general tasks (Vicuna Bench, MT Bench), retrieval-augmented generation (Neural-Bridge-RAG), and mathematical reasoning (MATH500, AMC23, GPQADiamond) - we demonstrate significant improvements in both effectiveness and efficiency compared to existing approaches.
@@ 2.1 Parallelization via Prompt Engineering
This category encompasses methods that leverage carefully crafted prompts to achieve parallel generation. SoT (Ning et al., 2023) adopts a two-phase approach: first generating skeletal outlines (e.g., bullet points), followed by concurrent branch generation through batch decoding or API calls. This introduces substantial overhead from KV-cache reinitialization during phase transitions. Additionally, this approach suffers from increased memory overhead due to batch processing and compromised response quality from constrained prompt templates.
PDOS (Yu, 2025) advances this approach by implementing internal parallel decoding through specialized masks and logits processors, eliminating the need for batching and multi-stage prompting. Nevertheless, as a prompt-based solution, it encounters inherent limitations: the model’s parallel capabilities remain underutilized, and the necessity for content re-prefilling during mode transitions diminishes efficiency gains.
@@ 2.2 Architecture-Modified Parallelization
  • Insufficient Intrinsic Parallelism Exploitation. APAR employs predefined rules and regular expressions to parallelize existing sequential responses, limiting its applicability to data with obvious formatting while struggling with mathematical and code content. PASTA parallelizes sequential responses through model rewriting but lacks further validation of independence and completeness between parallel branches. While achieving parallelism, these approaches compromise output quality or require non-trivial retraining costs.
In contrast, our ASPD framework better harnesses intrinsic parallelism through a non-invasive data transformation pipeline. By introducing branch-invisible masks and shared position encodings, we enable lossless and seamless transitions between serial and parallel decoding modes.
2.3 Orthogonal Acceleration Techniques
@@ 3.3 Hybrid Decoding Engine
During inference, when the model determines that the current response can be parallelized, it generates several parallel titles, each enclosed by <Title> and </Title>. After all branch titles have been generated, the model outputs a <Para> token, turning the engine into parallel decoding mode. For each parallel branch $i$, we set the prefix ”<Branch>Tii: ” to help the model identify which branch to generate, where Tii is the corresponding title for branch $i$. A branch is considered complete when the </Branch> token is reached. Once all branches are finished, a </Para> token is appended to the output sequence, switching the engine back to serial decoding mode.
The hybrid engine implements the branch-independent mask and synchronized position-id mechanism proposed in Section 3.2 to support parallel decoding. Each parallel token attends only to tokens from both the main branch and its respective branch, with sequentially increasing position-ids. This allows each branch to maintain the same generation pattern as native autoregressive models from its own perspective. After all parallel branches complete decoding, sequential decoding resumes on the main branch, where the position-id of the first token reflects its actual position in the complete sequence. Our hybrid decoding engine supports iterative serial-parallel switching for optimal efficiency.
@@ 4 Experiments
Building on our analysis of parallelization potential (Figure 1), our method effectively leverages intrinsic parallelism while maintaining generalizability across diverse scenarios (Figure 6). Results on Vicuna(Chiang et al., 2023), our Internal MRC, and RAG [^1] https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered benchmark demonstrate optimal parallelism utilization. The Mathematical Reasoning(Hendrycks et al., 2021) benchmark exhibits the highest parallelization constraints among evaluated tasks, with 88.4% PPD (+4.2%), 33.3% DP (+2.7%), and 2.6 ABN (-0.1), which exhibits strong alignment with the statistical patterns shown in Figure 1. These quantitative metrics provide compelling evidence that our approach successfully harnesses the inherent parallelization capabilities of LLMs while preserving the fundamental characteristics of native autoregressive generation.
@@ 4.1 Experimental Setup
Dataset. Following (Liu et al., 2024), we use ShareGPT Vicuna dataset [^1] https://huggingface.co/datasets/zwhe99/amc23 as our training data. This dataset encompasses instructions across various scenarios including STEM, roleplay, reasoning, and extraction tasks. Additionally, we utilize our internal Multi-turn Reading Comprehension (MRC) dataset for multi-dimensional validation. All data processing is conducted using our proposed non-invasive parallelization transformation pipeline to obtain parallelized data. For fair comparison, we create corresponding sequential data by removing special parallel tokens.
Evaluation. We conduct comprehensive evaluations using established benchmarks including Vicuna Bench (Chiang et al., 2023) and MT Bench (Zheng et al., 2023), following the evaluation protocol of (Liu et al., 2024). Furthermore, to evaluate models’ generalization capability , we introduce an out-of-domain benchmark, namely the RAG Bench. This benchmark consists of the first 200 questions with corresponding context sampled from rag-dataset-12000222https://huggingface.co/datasets/neural-bridge/rag-dataset-12000. Additionally, we evaluate on Internal Logical Reasoning (LR) and AI Search datasets. Our evaluation framework encompasses two primary dimensions: model effectiveness and computational efficiency.
  • Effectiveness Metrics. For public benchmarks, we utilize the LLM-as-judge evaluation framework proposed by (Zheng et al., 2023) to quantify response quality, maintaining methodological consistency with (Liu et al., 2024). All evaluations are conducted using Qwen3-235B-A22B (Yang et al., 2025) to ensure balanced assessment. For Internal datasets, we implement human evaluation protocols to mitigate potential model-based assessment biases.
@@
Table 1: Performance Comparison on MT Bench and Vicuna Bench. The bold and underlined values denote the best and second-best results, respectively. SoT represents evaluation results using official code on corresponding benchmarks, V-APAR shows results using open-source parameters, and V-APAR* corresponds to our fine-tuned APAR model using query-aligned data augmented by Qwen3-235B-A22B (Yang et al., 2025). Ori denotes the original model, while Seq refers to sequential fine-tuning without parallel tokens. Models evaluated include Vicuna-1.3-7B (V) and Qwen2.5-7B-Instruct (Q).
@@ 4.2 Main Results
Figure 7: Token Speed Analysis on Different Benchmarks. Our method achieves the best speedup performance on MT Bench and Vicuna Bench. For specific tasks, the speedup ratios reach 1.68-3.10x in scenarios like Generic, Writing, and CS.
@@
4.3 Generalization Across Different Datasets
To evaluate the generalization capability of parallel processing across different datasets, we conducted training on our internal MRC dataset and tested generalization performance on MRC, LR, and AI Search datasets, using Qwen2.5-7B-Instruct as our base model. Specifically, MRC is designed to evaluate a wide range of reading comprehension skills, including various question types such as single passage understanding, multi-information extraction, refusal to answer, etc. Logical Reasoning (LR) contains problems requiring step-by-step deductive reasoning, while AI Search resembles RAG scenarios where responses are generated by analyzing multi-source search results, typically presented in an itemized format.
As shown in Table 3, on the homogeneous MRC dataset, ASPD achieves a 1.35x speedup in TPS compared to the original model, while maintaining performance within 1% of the fine-tuned sequential model. Cross-validation using LR and AI Search datasets demonstrates acceleration ratios of 1.15x and 1.45x respectively.
@@ 4.3 Generalization Across Different Datasets
While previous approaches like APAR excluded mathematical and coding tasks from parallel data processing, we extend our investigation into these complex reasoning domains. Our evaluation includes a diverse range of benchmarks, including GPQA(Rein et al., 2024), Math500(Hendrycks et al., 2021), and competition-level mathematics such as AMC2311footnotemark: 1, AIME24222https://huggingface.co/datasets/HuggingFaceH4/aime_2024, and AIME25333https://huggingface.co/datasets/yentinglin/aime_2025. For training data, we utilize OpenR1-Math-220K [^4] https://huggingface.co/datasets/open-r1/OpenR1-Math-220k (Hugging Face, 2025), which was generated by prompting DeepSeek-R1(Guo et al., 2025) with NuminaMath1.5 (LI et al., 2024) as the query set. For optimal performance on these complex tasks, we utilized Qwen2.5-32B-Instruct(Team, 2024) as the backbone model. The model was trained for 9 epochs with a global batch size of 88, using a 12k context window and a learning rate of 1e-5.
@@
Parallelism Variation Across Difficulty Levels. As shown in Table 5, the DP significantly decreases from 33.30% in MATH500 to 8.6% in AIME2025, with TPS decreasing from 1.17x to 1.08x. This decline in parallelism indicates a strong correlation between performance degradation and difficulty level. Our detailed analysis of MATH500 in Figure 9(a) clearly shows that parallel efficiency metrics decrease as problem difficulty increases. This inverse relationship is consistently observed across various mathematical benchmarks in Figure 9(b), where higher difficulty levels corresponds to lower DP and TPS values. Notably, the GPQA dataset presents an interesting exception - its multiple-choice format naturally enables more parallel processing opportunities, leading to higher parallelization metrics despite comparable difficulty levels with other benchmarks.

### v2

@@ Abstract
The increasing scale and complexity of large language models (LLMs) pose significant inference latency challenges, primarily due to their autoregressive decoding paradigm characterized by the sequential nature of next-token prediction. By re-examining the outputs of autoregressive models, we observed that some segments exhibit parallelizable structures, which we term intrinsic parallelism. Decoding each parallelizable branch simultaneously (i.e. parallel decoding) can significantly improve the overall inference speed of LLMs. In this paper, we propose an Adaptive Serial-Parallel Decoding (ASPD), which addresses two core challenges: automated construction of parallelizable data and efficient parallel decoding mechanism. More specifically, we introduce a non-invasive pipeline that automatically extracts and validates parallelizable structures from the responses of autoregressive models. To empower efficient adaptive serial-parallel decoding, we implement a Hybrid Decoding Engine which enables seamless transitions between serial and parallel decoding modes while maintaining a reusable KV cache, maximizing computational efficiency. Extensive evaluations across General Tasks, Retrieval-Augmented Generation and Mathematical Reasoning demonstrate that ASPD achieves unprecedented performance in both effectiveness and efficiency. Notably, on Vicuna Bench, our method achieves up to 3.19x speedup (1.85x on average) while maintaining response quality within 1% difference compared to autoregressive models, realizing significant acceleration without compromising generation quality. Our framework sets a groundbreaking benchmark for efficient LLM parallel inference, paving the way for its deployment in latency-sensitive applications such as AI-powered customer service bots and answer retrieval engines.
@@ 1 Introduction
Figure 1: Overview of Data Intrinsic and Model Internal Parallelism. Proportion of Parallel Data (PPD). The ratio of parallelizable samples in the benchmark; Degree of Parallelism (DP). The proportion of parallel tokens among all tokens in the output; Average Branch Number (ABN). The average number of parallel branches in the output.
Through extensive analysis of LLM outputs, we have made a crucial observation: despite their sequential generation pattern, a substantial portion of the model’s outputs exhibits inherent parallelism, which can be harnessed for concurrent processing. As shown in Figure 1, across various scenarios including general [^1] https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered, retrieval-augmented-generation [^2] https://huggingface.co/datasets/neural-bridge/rag-dataset-12000, mathematical reasoning [^3] https://huggingface.co/datasets/open-r1/OpenR1-Math-220k , and our internal machine reading comprehension benchmarks, model responses consistently reveal substantial potential for parallelization. Within these parallelizable components, different scenarios exhibit varying levels of parallelism, ranging from 30.6% (Math) to 70.2% (General). By harnessing these inherently parallelizable responses for concurrent output, the model’s response speed can be significantly enhanced.
@@
Recent attempts to address these challenges have yielded promising but limited solutions. SoT (Ning et al., 2023) introduced a two-stage approach using response skeletons and parallel generation through external APIs, but this method incurs additional latency due to KV-cache limitations and increased memory overhead with batch processing. APAR (Liu et al., 2024) proposed discarding parallel branch’s KV-caches during final generation, which, while addressing position encoding issues, compromises response quality. PASTA (Jin et al., 2025) explored asynchronous parallel decoding with pre-allocated position ranges, but struggles with position encoding mismatches when actual generation lengths deviate from predictions.
@@
1) Non-invasive parallel data transformation pipeline. We develop an innovative pipeline that automatically discovers and extracts inherent parallelizable structures from autoregressive model responses, which identifies semantically independent components that can be processed concurrently while preserving the response’s original style. This enables us to build high-quality parallel training corpora automatically without altering the response probability distribution.
2) Internal parallelization module. Our framework introduces a custom branch-invisible mask that enables efficient parallel processing without the overhead of batching or threading, maintaining consistency with traditional autoregressive decoding patterns.
3) Hybrid Decoding Engine supports efficient and cyclic serial-parallel decoding. We propose novel parallel positional encoding and training techniques that enable zero-overhead integration of parallel branches and support multiple rounds of parallel decoding.
4) Comprehensive evaluation achieving an optimal balance between effectiveness and efficiency. Through extensive evaluation across diverse benchmarks - including general tasks (Vicuna Bench, MT Bench), retrieval-augmented generation (Neural-Bridge-RAG), and mathematical reasoning (MATH500, AMC23, GPQADiamond, AIME2024, AIME2025) - we demonstrate significant improvements in both effectiveness and efficiency compared to existing approaches.
@@ 2.2 Parallelization via Prompt Engineering
This category encompasses methods that leverage carefully crafted prompts to achieve parallel generation. SoT (Ning et al., 2023) adopts a two-phase approach: first generating skeletal outlines (e.g., bullet points), followed by concurrent branch generation through batch decoding or API calls. This introduces substantial overhead from KV-cache reinitialization during phase transitions. Additionally, this approach suffers from increased memory overhead due to batch processing and compromised response quality from constrained prompt templates. PDOS (Yu, 2025) advances this approach by implementing internal parallel decoding through specialized masks and logits processors, eliminating the need for batching and multi-stage prompting. Nevertheless, as a prompt-based solution, it encounters inherent limitations: the model’s parallel capabilities remain underutilized, and the necessity for content re-prefilling during mode transitions diminishes efficiency gains.
@@ 2.3 Architecture-Modified Parallelization
  • Insufficient Intrinsic Parallelism Exploitation. APAR employs predefined rules and regular expressions to parallelize existing sequential responses, limiting its applicability to data with obvious formatting while struggling with mathematical and code content. PASTA parallelizes sequential responses through model rewriting but lacks further validation of independence and completeness between parallel branches.
Our concurrent work Multiverse (Yang et al., 2025a), focusing on the Mathematical Reasoning task, also explores parallelism in large language model generation, implementing parallel branch generation through SGLang(Zheng et al., 2024) and leveraging Radix Attention for KV cache reuse. In contrast, our approach performs decoding within a single sequence, providing inherent continuity of the KV cache to directly reuse.
The proposed ASPD framework better harnesses intrinsic parallelism through a non-invasive data transformation pipeline. By introducing branch-invisible masks and shared position encodings, we enable lossless and seamless transitions between serial and parallel decoding modes. We systematically investigate diverse model architectures and inference paradigms , demonstrating significant acceleration effects without quality degradation across multiple domains, particularly in General Tasks and Retrieval-Augmented Generation scenarios.
@@ 3.3 Hybrid Decoding Engine
During inference, when the model determines that the current response can be parallelized, it generates several parallel titles, each enclosed by <Title> and </Title>. After all branch titles are generated, the model outputs a <Para> token, turning the engine into parallel decoding mode. For each parallel branch $i$, we set the prefix ”<Branch>Tii: ” to help the model identify which branch to generate, where Tii is the corresponding title for branch $i$. A branch is considered complete when the </Branch> token is reached. Once all branches are finished, a </Para> token is appended to the output sequence, switching the engine back to serial decoding mode.
The hybrid engine implements the branch-independent parallel mask and synchronized position-id mechanism proposed in Section 3.2 to support parallel decoding. Each parallel token attends only to tokens from both the main branch and its respective branch, with sequentially increasing position-ids. This allows each branch to maintain the same generation pattern as native autoregressive models from its own perspective. After all parallel branches complete decoding, sequential decoding resumes on the main branch, where the position ID of the first token reflects its actual position in the complete sequence. Our hybrid decoding engine supports iterative serial-parallel switching for optimal efficiency.
@@ 4 Experiments
Building on our analysis of parallelization potential (Figure 1), our method effectively leverages intrinsic parallelism while maintaining generalizability across diverse scenarios (Figure 6). Results on Vicuna(Chiang et al., 2023), our Internal MRC, and RAG [^1] https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered benchmark demonstrate optimal parallelism utilization. The Mathematical Reasoning (MATH500)(Hendrycks et al., 2021) benchmark exhibits the highest parallelization constraints among evaluated tasks, with 88.4% PPD (+4.2%), 33.3% DP (+2.7%), and 2.6 ABN (-0.1), which exhibits strong alignment with the statistical patterns shown in Figure 1. These quantitative metrics provide compelling evidence that our approach successfully harnesses the inherent parallelization capabilities of LLMs while preserving the fundamental characteristics of native autoregressive generation.
@@ 4.1 Experimental Setup
Dataset. Following (Liu et al., 2024), we use ShareGPT Vicuna dataset [^1] https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered as our training data. This dataset encompasses instructions across various scenarios including STEM, roleplay, reasoning, and extraction tasks. Additionally, we utilize our internal Machine Reading Comprehension (MRC) dataset for multi-dimensional validation. All data processing is conducted using our proposed non-invasive parallelization transformation pipeline to obtain parallelized data. For fair comparison, we create corresponding sequential data by removing special parallel tokens.
Evaluation. We conduct comprehensive evaluations using established benchmarks including Vicuna Bench (Chiang et al., 2023) and MT Bench (Zheng et al., 2023), following the evaluation protocol of (Liu et al., 2024). Furthermore, to evaluate models’ generalization capability , we introduce an out-of-domain benchmark, namely the RAG Bench. This benchmark consists of the first 200 questions with corresponding context sampled from rag-dataset-12000222https://huggingface.co/datasets/neural-bridge/rag-dataset-12000. Additionally, we evaluate on Internal Logical Reasoning (LR) and AI Search benchmarks. Our evaluation framework encompasses two primary dimensions: model effectiveness and computational efficiency.
  • Effectiveness Metrics. For public benchmarks, we utilize the LLM-as-judge evaluation framework proposed by (Zheng et al., 2023) to quantify response quality, maintaining methodological consistency with (Liu et al., 2024). All evaluations are conducted using Qwen3-235B-A22B (Yang et al., 2025b) to ensure balanced assessment. For Internal benchmarks, we implement human evaluation protocols to mitigate potential model-based assessment biases.
@@
Table 1: Performance Comparison on MT Bench and Vicuna Bench. The bold and underlined values denote the best and second-best results, respectively. SoT represents evaluation results using official code on corresponding benchmarks, V-APAR shows results using open-source parameters, and V-APAR* corresponds to our fine-tuned APAR model using query-aligned data augmented by Qwen3-235B-A22B (Yang et al., 2025b). Ori denotes the original model, while Seq refers to sequential fine-tuning without parallel tokens. Models evaluated include Vicuna-1.3-7B (V) and Qwen2.5-7B-Instruct (Q). CS is short for Common-Sense on Vicuna Bench.
@@ 4.2 Main Results
Figure 7: Token Speed Analysis on Different Benchmarks. Our method achieves the best speedup performance on MT Bench and Vicuna Bench. For specific tasks, the speedup ratios reach 1.68-3.10x in scenarios like Generic, Writing, and Common-Sense (CS).
@@
4.3 Generalization Across Different Benchmarks
To evaluate the generalization capability of parallel processing across different benchmarks, we conducted training on our internal MRC benchmark and tested generalization performance on MRC, LR, and AI Search benchmarks, using Qwen2.5-7B-Instruct as our base model. Specifically, MRC is designed to evaluate a wide range of reading comprehension skills, including various question types such as single passage understanding, multi-information extraction, refusal to answer, etc. Logical Reasoning (LR) contains problems requiring step-by-step deductive reasoning, while AI Search resembles RAG scenarios where responses are generated by analyzing multi-source search results, typically presented in an itemized format.
As shown in Table 3, on the homogeneous MRC benchmark, ASPD achieves a 1.35x speedup in TPS compared to the original model, while maintaining performance within 1% of the fine-tuned sequential model. Cross-validation using LR and AI Search benchmarks demonstrates acceleration ratios of 1.15x and 1.45x respectively.
@@ 4.3 Generalization Across Different Benchmarks
While previous approaches like APAR excluded mathematical and coding tasks from parallel data processing, we observe that our concurrent research work Multiverse(Yang et al., 2025a) has also directed attention toward mathematical problem parallelization. To thoroughly investigate this domain, we extended our model architecture to Qwen2.5-32B-Instruct(Team, 2024) and conducted comprehensive validation on these complex reasoning domains. Our evaluation includes a diverse range of benchmarks, including GPQA(Rein et al., 2024), MATH500(Hendrycks et al., 2021), and competition-level mathematics such as AMC23111https://huggingface.co/datasets/zwhe99/amc23, AIME24222https://huggingface.co/datasets/HuggingFaceH4/aime_2024, and AIME25333https://huggingface.co/datasets/yentinglin/aime_2025. For training data, we utilize OpenR1-Math-220K [^4] https://huggingface.co/datasets/open-r1/OpenR1-Math-220k (Hugging Face, 2025), which was generated by prompting DeepSeek-R1(Guo et al., 2025) with NuminaMath1.5 (LI et al., 2024) as the query set. The model was trained for 9 epochs with a global batch size of 88, using a 12k context window and a learning rate of 1e-5.
@@
Parallelism Variation Across Difficulty Levels. As shown in Table 5, the DP significantly decreases from 33.30% in MATH500 to 8.6% in AIME2025, with TPS decreasing from 1.17x to 1.08x. This decline in parallelism indicates a strong correlation between performance degradation and difficulty level. Our detailed analysis of MATH500 in Figure 9(a) clearly shows that parallel efficiency metrics decrease as problem difficulty increases. This inverse relationship is consistently observed across various mathematical benchmarks in Figure 9(b), where higher difficulty levels corresponds to lower DP and TPS values. Notably, the GPQA benchmark presents an interesting exception - its multiple-choice format naturally enables more parallel processing opportunities, leading to higher parallelization metrics despite comparable difficulty levels with other benchmarks.